OpenStudy (anonymous):

I am still confused on these two questions, can anyone help? Which of these are true or false? 1. sin(x)+cos(x) / sin(x) = cos(x) 2. 2sin(x) = sin(x) + sin(x)

OpenStudy (danjs):

\[\sin(x) + \frac{ \cos(x) }{ \sin(x) } = \cos(x)\] that is the way you typed it, or are both terms divided by sin(x) on the left side

OpenStudy (eric_d):

I think he meant like this..